# A Deep Implicit-Explicit Minimizing Movement Method for Option Pricing in Jump-Diffusion Models

### v1

@@ Abstract.
We develop a novel deep learning approach for pricing European basket options written on assets that follow jump-diffusion dynamics. The option pricing problem is formulated as a partial integro-differential equation, which is approximated via a new implicit-explicit minimizing movement time-stepping approach, involving approximation by deep, residual-type Artificial Neural Networks (ANNs) for each time step. The integral operator is discretized via two different approaches: a) a sparse-grid Gauss–Hermite approximation following localised coordinate axes arising from singular value decompositions, and b) an ANN-based high-dimensional special-purpose quadrature rule. Crucially, the proposed ANN is constructed to ensure the asymptotic behavior of the solution for large values of the underlyings and also leads to consistent outputs with respect to a priori known qualitative properties of the solution. The performance and robustness with respect to the dimension of the methods are assessed in a series of numerical experiments involving the Merton jump-diffusion model.
@@
68T07, 68T07, 91G60, 91G20, 65M12
This research work was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “First Call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment grant” (Project Number: 2152). The authors also acknowledge the use of computational resources of the DelftBlue supercomputer, provided by the Delft High-Performance Computing Centre [8]. EHG also wishes to acknowledge the financial support of The Leverhulme Trust (grant number RPG-2021-238) and of EPSRC (grant number EP/W005840/1).
@@ 1. Introduction
The industry standard for pricing European single-asset options in Lévy and affine jump-diffusion models are transform methods, see e.g. Carr and Madan [6], Fang and Oosterlee [13] or Eberlein et al. [12]. These methods, however, suffer from the curse of dimensionality and cannot be used for the valuation of basket options in dimensions higher than 7 or 8; see Bayer et al. [4] for a very recent article in this direction.
An alternative and general method for pricing options in Lévy and affine models is to solve the associated second-order partial integro-differential equation (PIDE), where the initial (terminal) condition is determined by the payoff function (1.1) of the option. The PIDE is derived by the Fundamental Theorem of Asset Pricing and the Feynman–Kac lemma, which relates the discounted expectation of the payoff with an initial (terminal) value problem of the form described in the following section.
@@
A further computational challenge stems from the complexity of the integral operator due to the jumps in the asset price process. To that end, we devise and compare two different approaches to address the curse of dimensionality there: a) a sparse-grid Gauss–Hermite approximation following localised coordinate axes arising from singular value decompositions, and b) an ANN-based high-dimensional special-purpose quadrature rule.
@@ 2.3. Implicit-explicit minimizing movement method
We will first describe the method for the PIDE defined on $\mathbb{T}\times\Omega$ with $\mathbb{T}=(0,T]$ and $\Omega=[0,x_{\mathrm{max}}]^{d}$, $x_{\max}>0$. In the next sections, we will also discuss specific modelling issues related to option pricing problems, which are, in turn, incorporated in order to extend the solution domain of the method. For the moment, and in a consistent fashion to what follows below, we prescribe homogeneous Dirichlet boundary conditions on $\partial\Omega$.
@@ 2.4. Decomposition of the solution
In the above discussion, we assumed, for simplicity, homogeneous Dirichlet boundary conditions on the truncated domain boundary $\partial\Omega$. In the option pricing context, however, one expects non-zero values of $u$ as $|x|\to\infty$. Fortunately, the modelling asserts that these values of the solution $u$ as $|x|\to\infty$ are asymptotically known. In particular, the value of the option at time $t\geq 0$ can be expressed as
@@
Therefore, the decomposition (2.14) implies that $\lim_{\sum a_{i}x_{i}\rightarrow\infty}w(t,x)=0$, which justifies the use of homogeneous Dirichlet boundary conditions on the domain $\Omega=[0,x_{\max})^{d}$ for $x_{\max}$ large enough, if we solve for $w$ instead.
@@ 2.5. Domain truncation
$u(t,x^{\prime})\approx u(t,x)+\sum_{i=1}^{d}\alpha_{i}(x_{i}^{\prime}-x_{i}).$ (2.18)
@@ 3.1. Network architecture
$+\tau\int_{\Omega}\mathcal{E}[U(t_{k},x,\theta)]\,\mathrm{d}x+\tau\sum_{j=0}^{p-1}\gamma_{j}\int_{\Omega}f[U(t_{j(k)},x,\theta^{j(k)})]U(t_{k},x,\theta)\,\mathrm{d}x,$
@@ 3.2.1. Initialization
with $\delta_{ij}$ being the Kronecker delta. In other words, we impose $w^{0}(x;\theta^{0})=0$ strongly at $t=0$, resulting to the network weights becoming independent of the initial condition of the PIDE.
@@
$w^{0}(x;\theta^{0})\to\min_{\theta}\|w^{0}(x;\theta)-f^{0}(x)\|^{2}_{L^{2}([0,x_{r}]^{d})}.$
@@ 3.3.1. Sparse Gauss–Hermite quadrature
In the Merton model, the integral can be expressed as the expected value, with respect to a Gaussian measure, of the function $h(z)=u(x\mathrm{e}^{z})-u(x)$ multiplied by the Poisson parameter $\lambda$; for brevity, for the remainder of this discussion we suppress the dependence on $t$, and $u$ signifies any of the $U(t_{j(k)},x,\theta^{j(k)})$ for $j=0,\dots,p-1$. In addition, the amplitude of the random jumps is determined through a multivariate normal random variable, $Z\sim\mathcal{N}(\mu_{J},\Sigma_{J})$. Therefore, $\Sigma_{J}$ introduces directional features.
@@
$\int_{\mathbb{R}^{d}}h(z)\nu(\mathrm{d}z)$ $=\lambda\pi^{-d/2}\int_{\mathbb{R}^{d}}h(\mu+\sqrt{2}Vy)\exp(y^{\mathsf{T}}y)\mathrm{d}y$
$\approx\lambda\pi^{-d/2}\sum_{\mathbf{i}\in\Theta_{q}}h(\mu+\sqrt{2}Vy^{\mathbf{i}})W^{\mathbf{i}},$ (3.6)
@@ 3.3.2. Approximation of the integral using an ANN
Concluding, it order to estimate the solution $U^{k}$ of the PIDE at time $t^{k}$, we need to determine the values of the integral operator. In order to estimate these values, we use the supplementary ANN. Therefore, we seek the combined optimizer $(\theta^{k},\phi^{k})$ for the expanded discrete cost functional, which is the sum of Equations (3.4) and (3.9).
@@ 4. Numerical examples
In this section, we test the performance of the proposed methodology by pricing a European basket call option in the Merton model. More specifically, we consider a basket of $d$ equally contributing assets with moneyness values $x_{1},x_{2},\dots,x_{d}$. The model parameters are chosen to be the following:
$\sigma_{i}=0.5,\ \rho_{ij}=\delta_{ij}+0.5(1-\delta_{ij}),\ \ i,j\in\mathbb{I},\ t\in\mathbb{T}=[0,T],$
@@
The first numerical example concerns the valuation of a European basket call option consisting of $5$ correlated assets. The maturity is chosen to be in one year, i.e. $T=1$. Initially, we use the Gauss–Hermite quadrature to compute the integral operator. Results for both the implicit-explicit Euler and the BDF-2 schemes are provided. For the Euler scheme, we consider a time-step of $\tau=0.01$, while the BDF-2 scheme uses a larger time-step of $\tau=0.04$. We present the option prices for moneyness values in the interval $[0,3]$, considering the scenario where all assets share the same moneyness. Figure 4.1 corresponds to the implicit-explicit Euler scheme, and Figure 4.2 to the BDF-2 scheme. For comparison reasons, we also provide an extended quasi Monte Carlo estimation of the solution. We can observe that both methods work very well compared to the Monte Carlo simulation, and the difference to the Monte Carlo is only visible for deep in-the-money options, with moneyness level above $2$. Indeed, the absolute error is on the order of $10^{-3}$. The implicit-explicit Euler method appears slightly more accurate than the BDF-2 scheme, due to the smallest time-stepping. This comes at a computational cost, since it is approximately $1.8$ times slower.
@@
Next, we present the corresponding results using the second ANN instead of the Gauss–Hermite quadrature to estimate the integral operator. Figures 4.3 and 4.4 depict the recovered option prices for the implicit-explicit Euler and the BDF-2 schemes, respectively. Again, we compare the results using a greedy quasi Monte Carlo simulation. We can observe that the ANN computation of the integral operator works also very well, and the absolute error remains in the order of $10^{-3}$. The ANN method though results in a computation of basket options prices that is roughly $1.7$ times faster than the corresponding method with the Gauss–Hermite quadrature.
@@
Figure 4.5 illustrates a comparison of the two approaches employed for the computation of the integral operator in the case of the 5-asset European basket option for moneynesses in $[0,3]$ at $t=0.01$. We can see that the two methods yield similar approximations of the integral operator. Let us point out that the Gauss–Hermite quadrature tends to outperform the ANN for extreme moneyness values (small or large) due to its sampling-independent nature. More specifically, the quadrature captures the asymptotic behavior of the integral for large $x$, aligning with the expected jump sizes, i.e.: $\lim_{x\rightarrow\infty}I(x)=x\mathbb{E}[\mathrm{e}^{z}-1]$. On the other hand, the ANN-based approach tends to yield smoother approximations, particularly for commonly sampled moneyness levels.
@@
We further test our numerical methodology in a rather challenging high-dimensional option pricing scenario. In particular, we consider the case of a 15-asset European basket call option, as described earlier, with the maturity time being nine months, i.e. $T=0.75$. We employ time steps of $\tau=0.01$ for the implicit-explicit Euler method and $\tau=0.03$ for the BDF-2 scheme, respectively. Figure 4.6 presents the pricing results for the implicit-explicit Euler scheme, whereas Figure 4.7 provides the corresponding results for the BDF-2 scheme. Both methods are performing very well once again, and the important observation is that the absolute error remains in the order of $10^{-3}$, although the dimension of the problem is now three times higher. The BDF-2 scheme is now 1.7 times faster than the implicit-explicit Euler method, however the time-step is now smaller compared to the case with $5$ assets.
@@ 4.4. Hyperparameters and performance
Here we report the hyperparameters, the computational setup, and the computational running times of our implementations. The neural networks used consist of $2$ DGM layers with a width of $2^{6}$ neurons each. The weights of these networks are initialized using the Xavier initialization. The optimization is driven by the Adam optimizer using a fixed learning rate $\alpha=3\cdot 10^{-4}$. During the initialization phase ($t=0$), the neural networks (for the solution and for integration where applicable) are trained for a total of $2^{15}$ epochs, using $2^{15}$ sampled points at each epoch. A combination of uniform sampling for the moneyness and then sampling from the Dirichlet distribution is used during this step, providing a better initial start for the subsequent training steps. For the first time step ($t=\tau$), we perform $2^{14}$ and $2^{15}$ epochs for the 5-asset and 15-asset cases, respectively. Subsequently, $2^{12}$ (5-assets) and $2^{13}$ (15-assets) epochs are performed for the following time steps. At each iteration, a total of $2^{11}d$ points are sampled in +$\mathbb{R}^{d}_{+}$. We have used scrambled Sobol sequences, cf. Sobol’ [28], Owen [22], for the sampling, which provide low-discrepancy quasi-random samples.
We run our computational algorithms on a computational node of the DelftBlue supercomputer. In particular, we used 12 CPU cores (Intel XEON E5), 64GB RAM, and an NVIDIA Tesla V100S with 32GB RAM. The run times in the implementations used to produce the numerical experiments suggest that the proposed ANN methodology is comparable to, and typically outperforms, the quasi Monte Carlo method for the simulation of the moneyness curves. The advantage, however, of the ANN approach is that it produces a reliable global space-time solution to the PIDE problem, which would be extremely computationally demanding to achieve with Monte Carlo. The availability of the global space-time numerical solution can be used for further study, e.g., of the Greeks.

### v2

@@ Abstract.
We develop a novel deep learning approach for pricing European basket options written on assets that follow jump-diffusion dynamics. The option pricing problem is formulated as a partial integro-differential equation, which is approximated via a new implicit-explicit minimizing movement time-stepping approach, involving approximation by deep, residual-type Artificial Neural Networks (ANNs) for each time step. The integral operator is discretized via two different approaches: (a) a sparse-grid Gauss–Hermite approximation following localised coordinate axes arising from singular value decompositions, and (b) an ANN-based high-dimensional special-purpose quadrature rule. Crucially, the proposed ANN is constructed to ensure the appropriate asymptotic behavior of the solution for large values of the underlyings and also leads to consistent outputs with respect to a priori known qualitative properties of the solution. The performance and robustness with respect to the dimension of these methods are assessed in a series of numerical experiments involving the Merton jump-diffusion model, while a comparison with the deep Galerkin method and the deep BSDE solver with jumps further supports the merits of the proposed approach.
@@
68T07, 65C20, 91G60, 91G20, 65M12
This research work was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “First Call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment grant” (Project Number: 2152). The authors also acknowledge the use of computational resources of the DelftBlue supercomputer, provided by the Delft High-Performance Computing Centre [9]. EHG also wishes to acknowledge the financial support of The Leverhulme Trust (grant number RPG-2021-238) and of EPSRC (grant number EP/W005840/2).
@@ 1. Introduction
The industry standard for pricing European single-asset options in Lévy and affine jump-diffusion models are transform methods, such as Fourier or COS; see e.g. Carr and Madan [7], Fang and Oosterlee [14] or Eberlein et al. [13]. These methods, however, suffer from the curse of dimensionality and can typically not be applied for the valuation of basket options in dimensions higher than 8 or 10; see Bayer et al. [4, 5] for the state of the art in this direction.
An alternative and general method for pricing options in Lévy and affine models is to solve the associated second-order partial integro-differential equation (PIDE), where the initial (terminal) condition is determined by the payoff function (1.1) of the option. The PIDE is derived by the Fundamental Theorem of Asset Pricing and the Feynman–Kac lemma, which relates the discounted expectation of the payoff with an initial (terminal) value problem of the form described in the following section. Once again, finite difference and finite element discretizations for the solution of the option pricing PIDEs suffer from the curse of dimensionality and cannot be used in dimensions higher than 6 or 8; see e.g. Griebel and Hullmann [17], Hepperger [19] and Reichmann and Schwab [28].
@@
A further computational challenge stems from the complexity of the integral operator due to the jumps in the asset price process. To that end, we devise and compare two different approaches to address the curse of dimensionality there: (a) a sparse-grid Gauss–Hermite approximation following localised coordinate axes arising from singular value decompositions, and (b) an ANN-based high-dimensional special-purpose quadrature rule.
@@
In addition, we provide a basic comparison of the proposed method against the popular deep Galerkin method (DGM) by Sirignano and Spiliopoulos [31] and the deep BSDE solvers with jumps by Han et al. [18] and by Gnoatto et al. [16], showcasing the competitiveness of the proposed method.
@@ 2.3. Implicit-explicit minimizing movement method
We will first describe the method for the PIDE defined on $\mathbb{T}\times\Omega$ with $\mathbb{T}=(0,T]$ and $\Omega=[0,x_{\mathrm{max}}]^{d}$, $x_{\max}>0$. In the next sections, we will also discuss specific modelling issues related to option pricing problems, which are, in turn, incorporated in order to extend the solution domain of the method. Let us, for the moment, prescribe homogeneous Dirichlet boundary conditions on $\partial\Omega$; this will be revisited in the next subsection.
@@ 2.4. Decomposition of the solution
In the above discussion, we assumed, for simplicity, homogeneous Dirichlet boundary conditions on the domain of the truncated boundary $\partial\Omega$. In the option pricing context, however, one expects non-zero values of $u$ as $|x|\to\infty$. Fortunately, the modelling asserts that these values of the solution $u$ as $|x|\to\infty$ are asymptotically known. In particular, the value of the option at time $t\geq 0$ can be expressed as
@@
Therefore, the decomposition (2.14) implies that $\lim_{\sum a_{i}x_{i}\rightarrow\infty}w(t,x)=0$, which justifies the use of homogeneous Dirichlet boundary conditions on the domain $\Omega=[0,x_{\max}]^{d}$ for $x_{\max}$ large enough, if we solve for $w$ instead.
@@ 2.5. Domain truncation
$u(t,x^{\prime})\approx u(t,x)+\sum_{i=1}^{d}\alpha_{i}(x_{i}^{\prime}-x_{i}),$ (2.18)
@@ 3.1. Network architecture
$\quad+\tau\int_{\Omega}\mathcal{E}[U(t_{k},x,\theta)]\,\mathrm{d}x+\tau\sum_{j=0}^{p-1}\gamma_{j}\int_{\Omega}f[U(t_{j(k)},x,\theta^{j(k)})]U(t_{k},x,\theta)\,\mathrm{d}x,$
@@ 3.2.1. Initialization
with $\delta_{ij}$ being the Kronecker delta. In other words, we impose $w^{0}(x;\theta^{0})=0$ strongly at $t=0$, resulting in the network weights becoming independent of the initial condition of the PIDE.
@@
$w^{0}(x;\theta^{0})\to\min_{\theta}\big{\|}w^{0}(x;\theta)-f^{0}(x)\big{\|}^{2}_{L^{2}([0,x_{r}]^{d})}.$
@@ 3.3.1. Sparse Gauss–Hermite quadrature
In the Merton model, the integral can be expressed as the expected value, with respect to a Gaussian measure, of the function $h(z)=u(x\mathrm{e}^{z})-u(x)$ multiplied by the Poisson parameter $\lambda$; for brevity, for the remainder of this discussion we suppress the dependence on $t$, and $u$ signifies any of the $U(t_{j(k)},x,\theta^{j(k)})$ for $j=0,\dots,p-1$. In addition, the amplitude of the random jumps is determined by a multivariate normal random variable, $Z\sim\mathcal{N}(\mu_{J},\Sigma_{J})$. Therefore, $\Sigma_{J}$ introduces directional features.
@@
$\int_{\mathbb{R}^{d}}h(z)\nu(\mathrm{d}z)$ $=\lambda\pi^{-d/2}\int_{\mathbb{R}^{d}}h(\mu+\sqrt{2}By)\exp(-y^{\mathsf{T}}y)\mathrm{d}y$
$\approx\lambda\pi^{-d/2}\sum_{\mathbf{i}\in\Theta_{q}}h(\mu+\sqrt{2}By^{\mathbf{i}})W^{\mathbf{i}},$ (3.6)
@@ 3.3.2. Approximation of the integral using an ANN
Concluding, in order to estimate the solution $U^{k}$ of the PIDE at time $t^{k}$, we need to determine the values of the integral operator. To approximate these values, we employ the supplementary ANN.
At each time step, we begin by optimizing the parameters $\phi^{k}$ of the ANN for the integral by minimizing (3.9), followed by the independent optimization for $\theta^{k}$ based on (3.4). Fortunately, the training associated with (3.9) is computationally inexpensive, introducing no significant overhead to the overall scheme.
@@ 4. Numerical examples
We shall now test the performance of the proposed methodology by pricing a European basket call option in the Merton model. Also, we provide a simple comparison with two popular and successful machine learning-driven solvers. More specifically, we consider a basket of $d$ equally weighted assets with moneyness values $x_{1},x_{2},\dots,x_{d}$. The model parameters are chosen to be the following:
$\sigma_{i}=0.5,\ \rho_{ij}=\delta_{ij}+0.5(1-\delta_{ij}),\ \ i,j\in\mathbb{I},\ t\in\mathbb{T}=(0,T],$
@@
The first numerical example concerns the valuation of a European basket call option consisting of $5$ correlated assets. The maturity is chosen to be in one year, i.e. $T=1$. Initially, we use the Gauss–Hermite quadrature to compute the integral operator. Results for both the implicit-explicit Euler and the BDF-2 schemes are provided. For the Euler scheme, we consider a time-step of $\tau=0.01$, while the BDF-2 scheme uses a larger time-step of $\tau=0.04$. We present the option prices for moneyness values in the interval $[0,3]$, considering the scenario where all assets share the same moneyness. Figure 4.1 corresponds to the implicit-explicit Euler scheme, and Figure 4.2 to the BDF-2 scheme. For comparison reasons, we also provide an extended Quasi Monte Carlo (QMC) estimation of the solution. We can observe that both methods work very well compared to the QMC simulation, and the difference to the Quasi Monte Carlo is only visible for deep in-the-money options, with moneyness level above $2$. Indeed, the absolute error is on the order of $10^{-3}$. The implicit-explicit Euler method appears slightly more accurate than the BDF-2 scheme, due to the smallest time-stepping. This comes at a computational cost, since it is approximately $1.8$ times slower.
@@
Next, we present the corresponding results using the second ANN instead of the Gauss–Hermite quadrature to estimate the integral operator. Figures 4.3 and 4.4 depict the recovered option prices for the implicit-explicit Euler and the BDF-2 schemes, respectively. Again, we compare the results using a greedy Quasi Monte Carlo simulation. We can observe that the ANN computation of the integral operator also works very well, and the absolute error remains in the order of $10^{-3}$. The ANN method though results in a computation of basket options prices that is roughly $1.7$ times faster than the corresponding method with the Gauss–Hermite quadrature.
@@
Figure 4.5 illustrates a comparison of the two approaches employed for the computation of the integral operator in the case of the 5-asset European basket option for moneynesses in $[0,3]$ at $t=0.01$. We can see that the two methods yield similar approximations of the integral operator. Let us point out that the Gauss–Hermite quadrature tends to outperform the ANN for extreme moneyness values (small or large) due to its sampling-independent nature. More specifically, the quadrature captures the asymptotic behavior of the integral for large $x$, aligning with the expected jump sizes, i.e.: $\lim_{x\rightarrow\infty}I_{\varphi}(x)=x\mathbb{E}[\mathrm{e}^{z}-1]$. On the other hand, the ANN-based approach tends to yield smoother approximations, particularly for commonly sampled moneyness levels.
@@
We further test our numerical methodology in a rather challenging high-dimensional option pricing scenario. In particular, we consider the case of a 15-asset European basket call option, as described earlier, with the maturity time being one year, i.e. $T=1$. We employ time steps of $\tau=0.01$ for the implicit-explicit Euler method and $\tau=0.03$ for the BDF-2 scheme, respectively. Figure 4.6 presents the pricing results for the implicit-explicit Euler scheme, whereas Figure 4.7 provides the corresponding results for the BDF-2 scheme. Both methods are performing very well once again, and the important observation is that the absolute error remains relative small, although the dimension of the problem is now three times higher. The BDF-2 scheme is now 1.7 times faster than the implicit-explicit Euler method, however the time-step is now smaller compared to the case with $5$ assets.
@@ 4.4. Hyperparameters and performance
We report the hyperparameters, the computational setup, and the computational run times of the implementation. The neural networks used consist of $2$ DGM layers with a width of $2^{6}$ neurons each. The weights are initialized via the Xavier initialization. The optimization is driven by the Adam optimizer using a fixed learning rate $\alpha=3\cdot 10^{-4}$. During the initialization phase ($t=0$), the neural networks (for the solution and for integration where applicable) are trained for a total of $2^{15}$ epochs, using $2^{15}$ sampled points at each epoch. A combination of uniform sampling for the moneyness and then sampling from the Dirichlet distribution is used during this step, providing a better initial start for the subsequent training steps. For the first time step ($t=\tau$), we perform $2^{14}$ and $2^{15}$ epochs for the 5-asset and 15-asset cases, respectively. Subsequently, $2^{12}$ (5-assets) and $2^{13}$ (15-assets) epochs are performed for the following time steps. At each iteration, a total of $2^{12}d$ points are sampled in +$\mathbb{R}^{d}_{+}$. We have used scrambled Sobol sequences, cf. Sobol’ [32], Owen [25], for the sampling, which provide low-discrepancy quasi-random samples. The algorithms have been executed on a computational node of the DelftBlue supercomputer. In particular, we used 12 CPU cores (Intel XEON E5), 64GB RAM, and an NVIDIA Tesla V100S with 32GB RAM.
4.5. Basic comparison with other ML solvers
In an effort to position the method proposed above in terms of computational cost, as well as comment on its benefits, we now provide a basic comparison against two popular alternative methods: the deep BSDE solver with jumps by Gnoatto et al. [16] and the deep Galerkin method by Sirignano and Spiliopoulos [31]; in the latter method, we incorporate the ANN approximation of the integral operator implicitly.
For the Deep BSDE solver, we adapt the code provided by Gnoatto et al. [16] on GitHub [^1] https://github.com/AlessandroGnoatto/DeepBsdeSolverWithJumps, which prices European basket call options under a jump-diffusion model with uncorrelated assets. To ensure fair comparison, we also consider uncorrelated assets for the other two methods.
We run the pricing schemes for a European basket call option with $5$ underlying assets, using the parameters: $T=1,\ K=1,\ \sigma_{i}=0.5,\ r=0.05,\ \lambda=1,\ \mu_{J_{i}}=0,\ \sigma_{Ji}=0.5$. For the deep BSDE solver, we solve for the initial moneyness value of $x_{i}=1$, while in the other two methods it is possible to price the option for an array of initial values at once.
Due to library incompatibilities with NVIDIA drivers, we have been unable to execute a CUDA-accelerated version of the deep BSDE solver. To facilitate a fair computational time assessment we, thus, compare CPU runtimes for all methodologies. We expect that CUDA execution would likely result in speedups of 10 times or more for all methods in a similar fashion. Moreover, to ensure a robust Monte Carlo (MC) estimation, we run 10 million iterations, leading to a standard deviation of $\mathrm{std}=1.147\cdot 10^{-4}$. Thus, we feel it is appropriate to consider the Monte Carlo solution as the ground truth.
Before providing the results of the basic comparison, we highlight the key differences on how these methods approach the solution space. The MC method solves for a single time-space point, providing an estimate for a specific set of initial conditions at maturity. The deep BSDE solver provides solutions across multiple time steps but for fixed spatial values, i.e., fixed initial asset or moneyness values. In contrast, both DGM and the deep IMEX scheme, proposed in this work, provide solutions that can be evaluated across the entirety of a space-time domain. This broader coverage makes DGM and deep IMEX more versatile but, as expected, they are also more computationally intensive. For instance, a key advantage of being provided with a complete solutions across the space-time domain, allows for straightforward calculation of Greeks. The latter is more challenging in local solvers, such as MC or deep BSDE.
We note that both deep IMEX and DGM are implemented with the same network architecture given in Section 3.1. To ensure, however, comparable accuracy, we test two different configurations for the DGM method: one with an additional layer and another with two additional layers, compared to the deep IMEX architecture. The results, including option values, absolute errors against the “ground truth” provided by MC, and runtimes, are summarized in Table 4.1. This basic comparison shows that the proposed deep IMEX approach is competitive with both MC and deep BSDE methodologies in terms of error and runtime, (less than twice as expensive as MC and less then three times as expensive as deep BSDE, for similar accuracy) to produce a complete space-time solution. Moreover, deep IMEX appears to be more economical than DGM in terms of runtime and parameter cardinality.
MC Deep IMEX Deep BSDE DGM (+1 layer) DGM (+2 layers)
Value $0.1647$ $0.1633$ $0.1661$ $0.1509$ $0.1669$
Abs. Error - $1.41E{-3}$ $1.36E{-3}$ $1.38E{-2}$ $2.14E{-3}$
Runtime (min) $533$ $885$ $376$ $1330$ $1634$
Table 4.1. Comparison of pricing methods for a European basket call option of 5 assets under a jump-diffusion model.